OpenStudy (anonymous):

esimate the circumfernece of a circle with the radious of 4 in. MEDAL + FAN!!!

OpenStudy (freckles):

State the formula for the circumference of a circle?

OpenStudy (anonymous):

so first we would have to find half of four right then times it by it then times it by 3.14??

OpenStudy (freckles):

once you can do that just replace r with 4

OpenStudy (anonymous):

circumference = 2*pi*r so circumference =2*3.14*4 = 25.12

OpenStudy (anonymous):

r is the radius = 4

OpenStudy (freckles):

or I guess someone can just do it for you

OpenStudy (anonymous):

well i appreciate both of your help thank you and you both gte fans :D

OpenStudy (anonymous):

thank you @sourav_aich i gave you medal and fan